RawlsK29:

I am so confused, how in the world do i solve this...? -1/2 - 4x/5 = 1/2 - 2x and this one... 2x - 3x/2 = -x + 1/3

RawlsK29:

@Ultrilliam Sorry i keep asking you questions but you seem to know what your doing😂

jhonyy9:

how you make calcules with fractions ? wat is the first step ? do you know what mean ,,common denominator" ?

jhonyy9:

@RawlsK29 are you here ? i like help you

RawlsK29:

Sorry i was getting a drink, yes im here. I have no idea what to do first

jhonyy9:

1. the first one add to both sides 1/2 so what will get ?

RawlsK29:

umm... the -1/2 cancels out so its -4x/5 = 1/2 - 2x

jhonyy9:

to both sides so this mean that you add to right side too 1/2 - ok ?

RawlsK29:

Oh okay i forgot about that. Soo... -4x/5 = idk

RawlsK29:

Do i add 1/2 + 1/2?

jhonyy9:

how you assume two fracvtions with the same denominator ? 1/a + 1/a = ?

RawlsK29:

multiply them?

jhonyy9:

no - how you ve learn it ? 1/a + 1/a = (1+1)/a or not ?

jhonyy9:

bc. a is the common denominator

jhonyy9:

so in this way 1/2 +1/2 = ?

RawlsK29:

2/2 = 1?

jhonyy9:

exactly - so in this way on the right side how will be ?

jhonyy9:

courage please - this is easy

RawlsK29:

Easy for you, i suck at math. how to solve -4x/5 = 1 - 2x? Its probably wrong, i told you i suck at math.

jhonyy9:

ok i m here to help you

jhonyy9:

-4x ---- = 1- 2x 5

RawlsK29:

So i got it right?😧

jhonyy9:

yes just now you need continue it how elimine this denominator of 5 from the left side ?

RawlsK29:

Umm... dont i flip the fraction then multiply then -4x/5 times 5/-4x Not sure though...

jhonyy9:

attention please than you multiply both sides by 5 what will get ?

RawlsK29:

Im confused

jhonyy9:

try multiply both sides by 5

RawlsK29:

Okay, so -4x = 5 - 2x??

jhonyy9:

-2x you not multiplied it by 5

RawlsK29:

-10x

jhonyy9:

on the right side what will get ?

jhonyy9:

-4x = ???

RawlsK29:

-4x = 5 - 10x?

jhonyy9:

exactly - very nice

jhonyy9:

and now ?

RawlsK29:

-4x = 5 - 10x +10 +10 6x = 5 x = 5/6

RawlsK29:

Is that right?

jhonyy9:

yes but you mised the x next 10 - so you need adding to both sides 10x

RawlsK29:

Oh okay, thanks! BTW your like super smart if you get this bull

jhonyy9:

wait one moment

RawlsK29:

k

jhonyy9:

6x=5 is just your first equation

RawlsK29:

I need another one...?😠

jhonyy9:

there is again the second one - so this you need solve it in the same way too

RawlsK29:

Oh dont worry about that, ill work on that tomorrow. Thanks anyways!

jhonyy9:

ok. this mean that you ve understood it now clearly sure - yes ?

RawlsK29:

Yes, cause you were an amazing explainer😂😂

jhonyy9:

ok. thank you very much

jhonyy9:

was my pleasure and anytime good luck bye bye